# Phase-shifted remote photoplethysmography for estimating heart rate and blood pressure from facial video

### 09 Jan 2024 v1

@@ 3.2 Preprocessing and ground truth generation
The preprocessing step extracts normalized ROI regions from the input images to remove the redundant background information. The pretrained MTCNN  (Zhang et al., 2016a) model was utilized to detect facial regions, and an ROI was decided to include all facial regions within a short video clip. The ROI regions are cropped and resized into the size of 128$\times$128, and the ROI images were normalized into the range between 0 and 1 to reduce the effect of light conditions.
On the other hand, pseudo PPG signals were generated from ABP signals by conducting detrending, bandpass filtering, and normalization. Detrending is the process of computing moving average of the ABP signals to remove high frequency noise components, and filtering was conducted to extract physiologically meaningful information from a pre-defined frequency band. Following the previous literatureHu et al. (2022), we used the smaller window size of 150 for detrending and extracted the information within the frequency band between 0.5 hz and 3 hz. The filtered signals were normalized into the range between -2 and 2 to generate pseudo PPG signals. Additionally, ground truth HR was computed by detecting the frequency value corresponding to the maximum power spectrum density. Ground truth values for SBP and DBP were computed by averaging peak and valley values within a time interval.
@@ 3.3 Data augmentation
Biometric information such as heart rate follows a Gaussian distribution and requires augmentation of long-tailed data due to data imbalance. In this paper, we propose upsampling and downsampling of the time axis based on frame interpolation to augment the data, and the process is shown in Fig. 2. In the training dataset, the original video sequence and signals with a window length of $T$ contain heart rate information of $r$ BPM. To augment the bradycardia data, the $\frac{T}{2}$ window length original video with a heart rate of $r$ BPM is used, and interpolation of the intermediate frames increases the sequence length to $T$ and reduces the heart rate to about $\frac{r}{2}$ BPM. The algorithm used for frame interpolation is FILM-Net (Reda et al., 2022) with transfer learning parameters. In contrast, the augmentation of tachycardia data uses the $2T$ window length original video with a heart rate of $r$ BPM, and skips an even number of frames to generate the $T$ window length sequence data with the heart rate of $2r$ BPM. The augmentation of heart rate data by upsampling and downsampling of frames has been proposed in Yu et al. (2020). The difference with our proposed method is that we use a deep learning-based frame interpolation method for upsampling.
@@ 3.4 Dual remote photoplethysmography network (DRP-Net)
In this paper, we propose a dual remote photoplethysmography network (DRP-Net) to estimate phase-shifted rPPGs from image sequences, whose structure is shown in Fig. 3. It is important for the rPPG estimation algorithm to learn spatiotemporal features to capture the time-series pulse wave signal $\textbf{y}_{PPG}\in R^{T}$ in the facial image sequence $X\in R^{3\times T\times 128\times 128}$. DRP-Net is a 3 dimensional CNN model, consisting of a feature extractor, spatial and temporal attention modules, and an rPPG prediction heads. The feature extractor employs atrous convolution to learn sparse local features over a large receptive field, while reducing the spatial dimension and maintaining the temporal dimension. The feature extractor produces a spatiotemporal feature map $\textbf{m}\in R^{64\times T\times 4\times 4}$, and it consists of the 7 atrous convolutions and 3 max pooling layers. In addition, the intermediate feature map $\textbf{m}_{inter}\in R^{64\times T\times 32\times 32}$ from the feature extractor is utilized in the spatial and temporal attention modules in parallel. The spatial attention module consists of 3 dimensional convolution, max pooling, average pooling, and pointwise convolution for dimensionality reduction to obtain $\textbf{v}_{s}\in R^{1\times 1\times 4\times 4}$. The spatial attention vector $\boldsymbol{\alpha}_{s}\in[0,1]$ is computed by applying the sigmoid function to $\textbf{v}_{s}$, and is defined as follows.
$\boldsymbol{\alpha}_{s}=\frac{exp(\textbf{v}_{s})}{1+exp(\textbf{v}_{s})},$ (1)
where $\boldsymbol{\alpha}_{s}$ separate the regions that contain physiological information from the background regions, giving the skin regions higher weights. Similarly, we obtain $\textbf{v}_{t}\in R^{1\times t\times 1\times 1}$ from the temporal attention module, and the temporal attention vector $\boldsymbol{\alpha}_{t}\in[0,1]$ is defined as
$\boldsymbol{\alpha}_{t}=\frac{exp(\textbf{v}_{t})}{1+exp(\textbf{v}_{t})},$ (2)
where $\boldsymbol{\alpha}_{t}$ gives relatively low weight to motion noise at the frame level and optimizes the model to minimize its impact. The rPPG prediction head estimates phase-shifted facial rPPG $\hat{\textbf{y}}_{f}\in R^{t}$ and acral rPPG $\hat{\textbf{y}}_{a}\in R^{t}$, and the input $\boldsymbol{\beta}\in R^{64\times T\times 4\times 4}$ is defined as follows.
@@
where $\otimes$ denotes elementwise multiplication, which gives the weights of temporal and spatial attention vectors $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\alpha}_{s}$ to spatiotemporal feature map m.
@@ 3.5 Bounded blood pressure network (BBP-Net)
To estimate SBP and DBP from facial rPPG $\hat{\textbf{y}}_{f}$ and acral rPPG $\hat{\textbf{y}}_{a}$, we designed a multi-scale feature fusion (MSF)-based bounded blood pressure network (BBP-Net), which is shown in Fig. 4. BBP-Net takes the $6\times T$ set of physiological signals as a stack of velocity plethysmograms (VPGs) and acceleration plethysmograms (APGs), which are the first and second derivatives of the facial and acral rPPG. The structure of BBP-Net is designed with a backbone network based on MSF blocks and a parallel BP prediction heads, and scaling sigmoid function. The MSF block performs depth-wise separable (DWS) convolution with kernel sizes of the 3 and 5 in parallel. Then it performs a linear operation and a channel-wise softmax operation to compute the weight vector. Every MSF blocks contain the residual connections. The BP prediction head includes the computation process to obtain the weight vector by employing the bottleneck attention module (BAM) (Park et al., 2018) and residual connections in parallel. We modify the sigmoid function into a scaling sigmoid function to bound the estimated blood pressure to the normal range, and it is defined as follows.
$\hat{y}_{BP}=BP_{min}+\frac{BP_{max}-BP_{min}}{1+exp({-z}+\tau)},\,where\,BP\in\left\{SBP,DBP\right\},$ (4)
where $\hat{y}_{BP}$ is the estimated blood pressure, and ${BP}_{max}$ and ${BP}_{min}$ are heuristic parameters that represent the range of reasonable blood pressure. $z$ and $\tau$ are the logit estimated from the BP prediction head and temperature, respectively. The structure of BBP-Net is based on the architecture proposed by  Hu et al. (2022), which is designed to estimate SBP and DBP from PPG, and has achieved state-of-the-art (SOTA) performance on the UCI-dataset. While previous studies that take a single PPG of size $3\times T$ as input, we utilize the phase-shifted information of the rPPG and employ a scaling sigmoid function.
@@ 3.6 Loss functions
Since the purpose of DRP-Net and BBP-Net is regression to estimate the biometrics of rPPG, heart rate, and blood pressure, we designed a loss function based on the mean squared error (MSE). The loss function $L_{HR}$ of DRP-Net for heart rate estimation is denoted as
$L_{HR}=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{y}_{HR,k}-y_{HR,k}\right)^{2},$ (5)
where $\hat{y}_{HR,k}$ and $y_{HR,k}$ are the predicted and ground truth heart rate, respectively, and $n$ is the number of data samples. The loss function in the frequency domain $L_{frequency}$ is defined based on the power spectrum of the physiological signal via FFT and can be defined as follows.
$L_{frequency}=\frac{1}{n}\sum_{k=1}^{n}\left(FFT(\hat{\textbf{y}}_{k})-FFT(\textbf{y}_{PPG,k})\right)^{2},$ (6)
where $\hat{\textbf{y}}_{k}$ and $\textbf{y}_{PPG,k}$ denote the rPPG predicted by DRP-Net and the PPG obtained from the sensor, respectively. In the case of using only frequency domain losses, the range of predicted signal’s amplitude is relatively small, therefore we propose $L_{peakvalley}$ as a loss function for scaling.
$L_{peakvalley}=\frac{1}{n}\sum_{k=1}^{n}\left[\left\{\phi(\hat{\textbf{y}}_{k}^{P})-\phi(\textbf{y}_{PPG,k}^{P})\right\}^{2}+\left\{\phi(\hat{\textbf{y}}_{k}^{V})-\phi(\textbf{y}_{PPG,k}^{V})\right\}^{2}\right],$ (7)
where $P$ and $V$ denote the peak and valley values of the signal, respectively, and the average of these values is denoted by $\phi$. $L_{peakvalley}$ is the time domain loss function but it is independent of the phase of the predicted signal. In addition, we define a loss function $L_{temporal}$ in the time domain to train an acral rPPG that simulates the scale and phase of the PPG. It is defined as
$L_{temporal}=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{\textbf{y}}_{k}-\textbf{y}_{PPG,k}\right)^{2}.$ (8)
The total loss function to optimize the facial rPPG is $L_{facial}=\lambda_{1}L_{HR}+\lambda_{2}L_{frequency}+L_{peakvalley}$, which guides the model to distinguish pulse waves in the facial image sequence. On the other hand, acral rPPG utilizes $L_{acral}=\lambda_{1}L_{HR}+\lambda_{2}L_{frequency}+L_{peakvalley}+L_{temporal}$ to simulate the phase of the PPG. BBP-Net for blood pressure estimation employs Huber loss (Collins, 1976) to predict SBP and DBP which we define as $L_{BP}$. Additionally, acral rPPG scaled with SBP and DBP ranges is adjusted by minimizing $L_{ABP}$ to the ABP signal.
@@ 4 Experimental results
Experiments were conducted on a hardware environment including Intel Core i9-10940X CPU, 64 GB DDR4 RAM, and NVIDIA Geforce RTX 3090 Ti. Python and Pytorch were used for the implementation of the proposed algorithm. For experimental details, a time window of 150 (6 seconds), a batch size of 8, and a decay factor of 0.99 were used, and the learning rate of the Adam optimizer was set to 0.001 and 0.0001 for MMSE-HR and V4V, respectively. To evaluate the accuracy of estimated heart rate and blood pressure, we adopted the mean absolute error (MAE) and root mean squared error (RMSE). Moreover, the proposed method was evaluated using the Pearson correlation coefficient r.

### 06 Mar 2024 v2

@@ 3.2 Preprocessing and ground truth generation
The preprocessing step extracts normalized ROI regions from the input images to remove the redundant background information. The pretrained MTCNN  (Zhang et al., 2016a) model was utilized to detect facial regions, and an ROI was decided to include all facial regions within a short video clip. The ROI regions are cropped and resized into the size of 128$\times$128, and the ROI images were normalized into the range between 0 and 1 to reduce the effect of light conditions. The sequence of ROI images was sampled at 25 frames per second (FPS) and split into window length of 150 frames (6 seconds) following the previous method (Chen et al., 2023).
The datasets utilized in this research contain ABP signals which carry various physiological information. The ABP signal was sampled and split at the same rate as the FPS and window length of 6 seconds of a facial video. Pseudo PPG signals were generated by conducting detrending, bandpass filtering, and standardization of ABP signals. Detrending is the process of computing time-varying finite-impulse response of the pseudo PPG signals  (Tarvainen et al., 2002) to remove high frequency noise components. Filtering was conducted within a pre-defined frequency band between 0.5 Hz and 3.0 Hz, enabling the acquisition of a wider range of physiologically meaningful information compared to the previous method (Nowara et al., 2021). The Fast Fourier Transform is used to extract the heart rate, which is computed by detecting the frequency corresponding to the maximum amplitude of the power spectral density (PSD) from the pseudo PPG signal. Ground truth values for SBP and DBP were computed by averaging the detected peaks and valleys within a window.
@@ 3.3 Data augmentation
We propose temporal upsampling and downsampling based on frame interpolation as an augmentation method for bradycardia and tachycardia data, and the process is shown in Fig. 2. Generally, the normal heart rate ranges between 60 and 100 BPM (Guyton, 2006), and bradycardia and tachycardia refer to heart rates lower and faster than normal heart rate, respectively. In the training dataset, An ROI sequence of the length of $T$ corresponding to the heart rate $r$ is augmented to generate bradycardia and tachycardia data which correspond to the heart rate of $\frac{T}{2}$ BPM and $2r$ BPM. To augment the bradycardia data, the $\frac{T}{2}$ frames of original ROI sequence is used, and expand the length of the window to $T$ frames via a frame interpolation method. The duration of the cardiac cycle decreases inversely with the increase in the sampling rate of the ROI sequence. The algorithm used for frame interpolation is pretrained FILM-Net (Reda et al., 2022). In contrast, the augmentation of tachycardia data downsamples $2T$ frames of the original ROI sequence by a factor of 2. The augmentation methods of remote physiological measurement include frame sampling (Yu et al., 2020; Lokendra & Puneet, 2022), color jitter (Yue et al., 2023), ROI masking (Chen et al., 2023), and synthetic data (Perche et al., 2021). The datasets utilized in this paper mostly contain a normal range of heart rates. To augment bradycardia and tachycardia data, we have improved the frame sampling methods.
@@ 3.4 Dual remote photoplethysmography network (DRP-Net)
In this paper, we propose a dual remote photoplethysmography network (DRP-Net) to estimate phase-shifted rPPG signals from image sequences, whose structure is shown in Fig. 3. It is important for the rPPG estimation model to learn spatiotemporal features to capture the time-series pulse wave signal in the facial image sequence $\textbf{X}\in\mathbb{R}^{3\times T\times H\times W}$. $H$ and $W$ denote height and width, and we set both of these hyperparameters of 128 in our experiments. DRP-Net is a three-dimensional CNN model, consisting of a feature extractor, spatial and temporal attention modules, and rPPG prediction heads. The feature extractor employs atrous convolution to learn sparse local features over a large receptive field, while reducing the spatial dimension and maintaining the temporal dimension. The kernel size and dilation ratio for each convolution layer are denoted by $k$ and $d$ in Fig. 3. The feature extractor consists of the seven atrous convolutions and three max pooling layers and produces a spatiotemporal feature map $\textbf{m}\in\mathbb{R}^{64\times T\times 4\times 4}$. In addition, the intermediate feature map $\textbf{m}_{inter}\in\mathbb{R}^{64\times T\times 32\times 32}$ from the feature extractor is utilized in the spatial and temporal attention modules in parallel. The spatial attention module consists of three-dimensional convolution, max pooling, average pooling, and pointwise convolution for dimensionality reduction to obtain $\textbf{v}_{s}\in\mathbb{R}^{1\times 1\times 4\times 4}$. The spatial attention vector $\boldsymbol{\alpha}_{s}$ is computed by applying the sigmoid function $\sigma$ to $\textbf{v}_{s}$. $\boldsymbol{\alpha}_{s}$ gives a higher weight to skin regions to distinguish areas containing physiological information and irrelevant background. Similarly, we obtain $\textbf{v}_{t}\in\mathbb{R}^{1\times T\times 1\times 1}$ from the temporal attention module, and acquire the temporal attention vector $\boldsymbol{\alpha}_{t}$. $\boldsymbol{\alpha}_{t}$ gives relatively low weight to motion noise at the frame level and optimizes the model to minimize its impact. The rPPG prediction heads consist of Siamese structured CNNs and take $\boldsymbol{\beta}$ as an input feature vector, defined as follows.
@@
where $\otimes$ denotes elementwise multiplication, which broadcasts vectors to a common shape and gives the weights of temporal and spatial attention vectors $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\alpha}_{s}$ to spatiotemporal feature map m. Each of the rPPG prediction heads predicts a one-dimensional signal $\hat{\textbf{y}}\in\mathbb{R}^{T}$, representing phase-shifted facial rPPG signals $\hat{\textbf{y}}_{f}$ and acral rPPG signals $\hat{\textbf{y}}_{a}$, respectively.
@@ 3.5 Bounded blood pressure network (BBP-Net)
We designed a multi-scale feature fusion(MSF)-based bounded blood pressure network (BBP-Net) to estimate SBP and DBP from facial rPPG signal $\hat{\textbf{y}}_{f}$ and acral rPPG signal $\hat{\textbf{y}}_{a}$, as shown in Fig. 4. BBP-Net takes the physiological signals as a stack of phase-shifted rPPG signals, velocity plethysmography (VPG) and acceleration plethysmography (APG) signals. VPG and APG are the first and second derivatives of the facial and acral rPPG signals. The structure of BBP-Net is designed with a backbone network based on MSF blocks and parallel BP prediction heads, and scaled sigmoid function. $C_{in}$, $C_{out}$, and $C_{mid}$ denote the channels of the input, output, and middle layers of the MSF blocks and the BP prediction heads, respectively, while $T_{in}$ denotes the length of the input signals. The MSF block performs depth-wise separable convolution with kernel sizes of the 3 and 5 in parallel and uses Hardswish (Howard et al., 2019) as an activation function. Then it performs a linear operation and a channel-wise softmax operation to compute the weight vector. All of MSF blocks contain the residual connections. The BP prediction head includes the computation process to obtain the weight vector by employing the bottleneck attention module (BAM) (Park et al., 2018) and residual connections in parallel. We modify the sigmoid function into a scaled sigmoid function to bound the estimated blood pressure to the predefined range, and it is defined as follows.
$\widehat{BP}=BP_{min}+(BP_{max}-BP_{min})\sigma(z/\tau),\,$ (2)
where $\widehat{BP}$ is the estimated blood pressure, and ${BP}_{max}$ and ${BP}_{min}$ are heuristic parameters that represent the range of reasonable blood pressure. The symbol $BP$ consists of two types of blood pressure, $SBP$ and $DBP$, and the predefined bounded ranges of $\widehat{SBP}$ and $\widehat{DBP}$ are set to [155, 85] and [95, 45] mmHg, respectively. $\sigma$, $z$ and $\tau$ denote the sigmoid function, the logit estimated from the BP prediction head and temperature, respectively. We set temperature to 2 in our experiments. The structure of BBP-Net is based on the architecture proposed by  Hu et al. (2022), which is designed to estimate SBP and DBP from PPG signals. While previous studies that take a stack of PPG, VPG, APG signals as input, we utilize the phase-shifted information of the rPPG signals and employ a scaled sigmoid function.
@@ 3.6 Loss functions
The loss functions for training DRP-Net are $L_{freq}$ and $L_{HR}$ in the frequency domain, and $L_{time}$ and $L_{pv}$ in the time domain. The frequency domain loss $L_{freq}$ is designed based on the PSD of the physiological signals and defined as
$L_{freq}=\lVert P(\hat{\textbf{y}})-P(\textbf{y})\lVert_{2},$ (3)
where $\hat{\textbf{y}}$ and y denote the rPPG signals predicted by DRP-Net and the pseudo PPG signals obtained from the ABP signals, respectively. $P(\cdot)$ indicates the PSD of the signals, which produces a real-valued vector obtained after applying the fast Fourier transform. The frequency range of the vectors is between 0.5 Hz and 3 Hz, corresponding to heart rates between 30 and 180 BPM. Moreover, the heart rate corresponds to the frequency with the maximum amplitude in the PSD, indicating the distinct period in the physiological signal. The loss function of heart rate $L_{HR}$ is defined as follows,
$L_{HR}=\lVert\widehat{HR}-HR\rVert_{2},$ (4)
where $\widehat{HR}$ and $HR$ are the predicted and ground truth heart rate, respectively. In addition, we define the time domain loss $L_{time}$ to estimate the acral rPPG signal, which mimics the scale and phase of the pseudo PPG signal, with the following equation (5).
$L_{time}=\lVert\hat{\textbf{y}}-\textbf{y}\rVert_{2}.$ (5)
We propose an additional time domain loss $L_{pv}$, which is phase-independent and designed to align the scale of facial rPPG signals. To remove noise and improve smoothness, we performed standardization, detrending, and filtering on the rPPG signals. We extract all of temporal location where the first derivative of the signal changes from positive to negative to detect all peaks in the window. This process is defined as follows,
$\tilde{\textbf{D}}^{r}(s_{t})=\left\{s_{t}\mid(s_{t}-s_{t-1})(s_{t+1}-s_{t})<0\right\}\left\{\begin{matrix}r=p~{}~{}if~{}s_{t}>s_{t-1},\\ r=v~{}~{}if~{}s_{t}<s_{t-1,}\end{matrix}\right.$ (6)
where $\tilde{\textbf{D}}^{r}(\cdot)$, $s$, $t$, $p$, and $v$ denote set of detected peaks or valleys, input signal, sample of signal, peak and valley, respectively. The values lower than the average of detected peaks are removed to minimize the effect of the dicrotic notch in each pulse. The remaining peaks indicate a vector composed of the maximum value of each pulse. The loss is calculated as the average of the remaining peaks using the following equation (7), considering temporal discrepancy and difference of vector sizes between rPPG and pseudo PPG signals.
$D^{p}(s)=\left\{\Phi(\tilde{\textbf{D}}_{i}^{p}(s))\mid\tilde{\textbf{D}}_{i}^{p}(s)>\Phi(\tilde{\textbf{D}}^{p}(s))\right\},$ (7)
where $D^{p}(\cdot)$, $\Phi(\cdot)$, and $i$ denote peak value within the window, average, and $i$th sample of detected peaks, respectively. In the case of valleys, the standards for detection and removal are opposite to those for peaks, as defined as follows,
$D^{v}(s)=\left\{\Phi(\tilde{\textbf{D}}_{i}^{v}(s))\mid\tilde{\textbf{D}}_{i}^{v}(s)<\Phi(\tilde{\textbf{D}}^{v}(s))\right\}.$ (8)
$L_{pv}$ is defined based on the L2 norm of the peak valley matrix of the predicted rPPG and pseudo PPG signals, with the following equation,
$L_{pv}=\lVert\hat{\textbf{c}}-\textbf{c}\rVert_{2},~{}where~{}\hat{\textbf{c}}=[D^{p}(\hat{y}),D^{v}(\hat{y})],\textbf{c}=[D^{p}(y),D^{v}(y)],$ (9)
where $\hat{\textbf{c}}$ and c denote the peak valley matrix of predicted rPPG and pseudo PPG signals, respectively.
The total loss function to optimize the facial rPPG signals is defined as
$L_{facial}=\lambda_{1}L_{HR}+\lambda_{2}L_{freq}+L_{pv},$ (10)
which guides the model to extract pulse waves in the facial image sequence. On the other hand, optimization of acral rPPG signals utilizes the following equation,
$L_{acral}=\lambda_{1}L_{HR}+\lambda_{2}L_{freq}+L_{pv}+L_{time},$ (11)
to mimic the phase of the pseudo PPG signal which obtained from acral site. We experimented by setting the constant $\lambda_{1}$ and $\lambda_{2}$ to 0.0001 and 100, respectively.
We designed $L_{BP}$ and $L_{ABP}$ as loss functions of BBP-Net for blood pressure estimation. $L_{BP}$ employs Huber loss (Collins, 1976) to calculate the loss for the predicted SBP and DBP, defined by the following equation,
$L_{BP}=\left\{\begin{matrix}\frac{1}{2}(\widehat{BP}-BP)^{2}&if\left|(\widehat{BP}-BP)\right|<\delta,\\ \delta\left|(\widehat{BP}-BP)\right|-\frac{1}{2}\delta),&otherwise,\end{matrix}\right.$ (12)
where $\widehat{BP}$ and $BP$ denote for predicted and ground truth blood pressure and heuristic parameter $\delta$ which set to 1 in the experiment, determines the type of loss function for the error. The Huber loss is based on the L2 and L1 norm, applying the former for errors within a defined $\delta$ range for smoothness and differentiability, and the latter for larger errors to enhance robustness against outliers. In addition, we reconstructed the ABP signal in the time domain using the predicted physiological information, scaling $\hat{\textbf{y}}_{a}$ to the range of $\widehat{SBP}$ and $\widehat{DBP}$ as follows,
$\hat{\textbf{y}}_{s}=\frac{\hat{\textbf{y}}_{a}-\hat{y}_{min}}{\hat{y}_{max}-\hat{y}_{min}}(\widehat{SBP}-\widehat{DBP})+\widehat{DBP},$ (13)
where $\hat{y}_{max}$ and $\hat{y}_{min}$ are the maximum and minimum values of $\hat{\textbf{y}}_{a}$, and $\hat{\textbf{y}}_{s}$ denotes scaled acral rPPG signal. $L_{ABP}$ is defined based on the L2 norm of scaled acral rPPG signal $\hat{\textbf{y}}_{s}$ and the ABP signal Y, defined as follows,
$L_{ABP}=\lVert\hat{\textbf{y}}_{s}-\textbf{Y}\rVert_{2}.$ (14)
$L_{ABP}$ fine-tunes $\widehat{SBP}$ and $\widehat{DBP}$ to exist within the ranges of peak and valley of the ABP signal as they change from pulse to pulse.
@@ 4 Experimental results
Experiments were conducted on a hardware environment including Intel Core i9-10940X CPU, 64 GB DDR4 RAM, and NVIDIA Geforce RTX 3090 Ti. Python and Pytorch were used for the implementation of the proposed algorithm. For experimental details, a window of 150 (6 seconds), a batch size of 8, and a decay factor of 0.99 were used, and the learning rate of the Adam optimizer was set to 0.001 and 0.0001 for MMSE-HR and V4V, respectively. To evaluate the accuracy of estimated heart rate and blood pressure, we adopted the mean absolute error (MAE), root mean squared error (RMSE), Pearson correlation coefficient r.